# SeRL: Self-Play Reinforcement Learning for Large Language Models with Limited Data

### v1

@@ 4.2 Experimental Results
Our approach matches the performance of the method using extensive data with verifiable rewards. As shown in Tab. 1, after just the first round of unsupervised RL training on generated data, our model achieves performance comparable to RL-GT. After the second round, LLaMA-3.2-3B-Instruct even outperforms RL-GT across nearly all evaluation benchmarks. LLaMA-3.2-3B-Instruct continues to improve in the third round. After three rounds of iterative training, Qwen-2.5-7B-Instruct also outperforms RL-GT on the MATH-500, ASDiv, and TabMWP benchmarks.
Our method outperforms other multi-round iterative approaches. Across all iterations, our method consistently outperforms both SR-DPO and I-RPO on LLaMA-3.2-3B-Instruct. We observe that both SR-DPO and I-RPO exhibit noticeable model degradation over multiple iterations. For SR-DPO, it performs poorly on LLaMA-3.2-3B-Instruct, while showing slightly better results on Qwen-2.5-7B-Instruct. This suggests that model-based reward mechanisms may depend on model scale, as larger models are more likely to produce accurate reward estimates. To better understand the limited effectiveness of SR-DPO, we further analyze its reward estimation accuracy in Section 4.3. In contrast, I-RPO shows a decline in performance over multiple training rounds on both models. We hypothesize that this degradation results from the inclusion of a negative log-likelihood (NLL) loss term in its training objective, which may cause the model to overfit the chosen responses. As a consequence, this overoptimization reduces generalization and ultimately leads to performance drops.
Our method generalizes well to general reasoning tasks. To better evaluate generalization, we assess our method on the MMLU-Pro benchmark. As shown in Tab. 2, LLaMA-3.2-3B-Instruct achieves consistent gains over multiple iterations in certain categories, whereas Qwen-2.5-7B-Instruct shows minimal improvement. We attribute this to Qwen-2.5 already being extensively fine-tuned on MATH data, as further rule-based RL training on the MATH training set also fails to improve its MMLU-Pro performance. LLaMA-3.2-3B-Instruct shows consistent improvement in the STEM category, with additional gains in Humanities, eventually surpassing RL-GT. Performance in the Social and Other categories remains largely unchanged. This is likely because our training data is math-focused, enhancing logical reasoning skills that benefit STEM tasks. The improvement in Humanities is mainly due to the Law subcategory, where many questions involve logical reasoning. As the reasoning ability of the model improves, its performance on Law tasks also increases.
@@ 4.3 Ablation Study
The dual-end clipped difficulty filtering mechanism effectively prevents reward hacking. As shown in Fig. 2, when we remove the difficulty filtering, the model exhibits clear signs of reward hacking: the average reward continues to increase with training steps, but the accuracy on the MATH-500 test set drops sharply after around 100 steps.
@@
Figure 2: Training curve of LLaMA-3.2-3B-Instruct without the online filter strategy.
[…]
@@
Our Self-Rewarding Method is Effective. Self-rewarding methods are fundamentally designed to approximate the true reward. Their effectiveness can be evaluated by measuring how closely the estimated rewards align with the ground-truth rewards, as greater agreement indicates more accurate reward estimation. To quantify this, we calculate the cosine similarity between the estimated rewards and the ground-truth rewards, with higher similarity indicating greater accuracy in reward estimation. We sample 16 responses from the MATH-500 dataset and computed rewards for each using three methods: majority-vote-based reward, model-based reward and rule-based reward. The rule-based reward is regarded as the ground-truth reward, against which we compute the cosine similarity of other alternative reward estimates. As shown in Fig. 3, on both LLaMA-3.2-3B-Instruct and Qwen-2.5-7B-Instruct, the majority-vote-based reward demonstrates significantly higher alignment with the rule-based reward, indicating superior accuracy in reward estimation. In contrast, the low accuracy of the model-based reward on LLaMA-3.2-3B-Instruct likely accounts for the underperformance of SR-DPO on this model. Additionally, the model-based reward shows higher consistency with the rule-based reward on Qwen-2.5-7B-Instruct, which accounts for the relatively better performance of SR-DPO on Qwen-2.5-7B-Instruct.
@@
Figure 3: Cosine similarity between rule-based reward and majority-vote-based / model-based reward for each instruction on MATH500. Each point represents the similarity over 16 sampled responses per instruction. "sim(GT, Model-based)" refers to the cosine similarity between the rule-based Reward and the model-based Reward, while "sim(GT, Majority-vote-based)" measures the cosine similarity between the rule-based reward and our majority-vote-based reward. The dashed lines in the figure indicate the average values.
To further assess the effectiveness of our majority-vote-based reward, we conduct an ablation study shown in Tab. 3. In this setting, we remove the self-instruction mechanism and perform one round of majority-vote-based reward RL using the full MATH training set, denoted as RL-MV. Results show that this variant achieves performance comparable to RL-GT across all test datasets. Notably, LLaMA-3.2-3B-Instruct outperforms RL-GT on MATH-500, MATH-Hard, College Math, and TabMWP, while Qwen-2.5-7B-Instruct consistently surpasses RL-GT on nearly all benchmarks.
@@
Impact of RL Algorithms. To verify the robustness of our method across different RL algorithms, we evaluate it under various RL setups. RLOO (Ahmadian et al., 2024) and Reinforce++ (Hu et al., 2025) use $k_{1}$ KL divergence estimation, while GRPO (Shao et al., 2024) adopts $k_{3}$ KL divergence estimation. The formulations for both $k_{1}$ and $k_{3}$ KL divergence estimations are detailed in Appendix B. Since $k_{1}$ is integrated into the per-token reward term, we set a smaller KL coefficient for training RLOO and Reinforce++ to avoid strong KL regularization. All other hyperparameter settings remain the same. As shown in Tab. 4, all algorithms achieve comparable performance, demonstrating the generality of our approach. Among them, the Reinforce++ algorithm delivers the best overall results. The GRPO algorithm achieves relatively strong performance in the first iteration but shows continued improvement only on MATH-Hard, ASDiv, and TabMWP in subsequent rounds, while its performance declines on MATH-500 and College Math. Similarly, RLOO exhibits performance degradation on MATH-Hard and College Math after the third iteration. Training curves reveal that both RLOO and GRPO experience more rapid KL divergence growth compared to Reinforce++, suggesting greater deviation from the initial model. Such a large divergence can lead to over-optimization and subsequent performance degradation. Furthermore, due to the exponential term in the $k_{3}$ KL divergence estimation, GRPO occasionally exhibits abrupt spikes in KL values, potentially causing training instability. Consequently, Reinforce++ stands out as the most stable and robust algorithm across multiple iterations.
Table 4: Comparison of SeRL performance under different reinforcement learning algorithms. Bold indicates the best performance of each RL method across iterations, and underlined denotes the best result among all methods and iterations.
@@ 5 Related Work
Self-Instruction Methods. Recent studies show that RL greatly enhances the reasoning abilities of LLMs, but relies on high-quality instructions and verifiable labels, which are scarce in specified domains such as clinical diagnostics (Ullah et al., 2024; McDuff et al., 2025; Wang and Liu, 2025) and aerospace engineering (Liu et al., 2025a; Connolly, 2025; Yadav, 2024). This has led to growing interest in synthetic data generation. Prior work has leveraged expert models such as GPT-3.5 or GPT-4 to generate synthetic data (Yang et al., 2023; Goldie et al., 2025; Li et al., 2024b, a; Yuan et al., 2025b). While effective, these methods rely on powerful external models which is not consistently available and often incur high construction costs. To reduce reliance on external help, self-instruction frameworks such as Wang et al. (2023) bootstrap data from a small set of seed datasets using few-shot generation and apply heuristic filtering to ensure the quality and diversity of generated data. Many other studies (Zhao et al., 2024; GALLOY et al., 2025; Kim et al., 2025; Gao et al., 2024; Guo et al., 2024) have extended the self-instruction framework, focusing on improving data efficiency, diversity, and quality. Nonetheless, these approaches are restricted to offline SFT, where fixed data cannot adapt to the evolving ability of the model. In contrast, RL enables online data generation that better aligns with the current capability of the model. However, RL requires not only instructions but also verifiable labels and mechanisms to prevent reward hacking.
@@
Self-Rewarding Methods. Recent successes of RL in LLMs largely rely on large-scale instructions with verifiable rewards (Xiong et al., 2025; Jiao et al., 2025; Wen et al., 2024). However, labeled data typically requires substantial human effort and is limited in both efficiency and quality. As a more efficient alternative, self-rewarding methods that estimate rewards without relying on verifiable labels have gained increasing attention. Model-based self-rewarding methods like Bai et al. (2022) and Yuan et al. (2025a) use predefined rules to generate preference pairs for Direct Preference Optimization (DPO) training. Fränken et al. (2024) maximize mutual information between principles and responses, while Zhang et al. (2025b) is based on semantic entropy clustering. TTRL (Zuo et al., 2025), as a concurrent work, also adopts a majority-vote-based self-rewarding mechanism in an unsupervised RL setting. However, our approach differs in two key ways: (1) we generate synthetic training instructions from limited initial data via self-instruction, unlike test-time training of TTRL on fixed test sets; (2) our method supports multi-round iterative updates, enabling continual self-improvement, whereas TTRL performs training only once on a fixed test set.
@@ 6 Discussion
Ceiling of Self-Iteration. We begin by discussing the performance upper bound of LLMs within a single iteration. When performing RL training using only majority-vote-based reward, the maximum achievable Pass@1 on the training set is inherently limited by the Maj@K before training. If the LLM has strong generalization capabilities, the Pass@1 upper bound on an in-domain test set should also align with the Maj@N of the model on the test set before training. So, what is the upper bound of LLM capability across multiple iterations? If the Maj@N of the model continues to improve after each round of training, then each iteration defines a new, higher ceiling. In this case, the LLM is capable of approaching increasingly better performance bounds, suggesting the potential for unbounded, continual improvement through iterative self-training. However, as shown in Fig. 7, the Maj@16 performance of both LLaMA-3.2-3B-Instruct and Qwen-2.5-7B-Instruct on MATH-500 does not continue to improve with more iterations. This observation aligns with the analysis by Yue et al. (2025), which suggests that RL with verifiable reward primarily transforms Pass@K ability into Pass@1, while the upper bound of a model is constrained by its underlying capability. Nevertheless, our method remains valuable, as it enables efficient reasoning in data-scarce scenarios through unsupervised RL training with data augmentation.
@@ 7 Conclusion
In this work, we introduce SeRL, a framework designed to bootstrap LLM training from limited initial data. SeRL consists of two key components: a self-instruction module and a self-rewarding module. The self-instruction module expands the initial data by generating new instructions, with robust and effective online filtering applied to ensure quality, diversity, and appropriate difficulty. The self-rewarding module uses a majority-vote strategy to estimate rewards for generated responses, removing the need for external labels. Based on the generated data, SeRL performs standard RL training in an iterative, self-improving manner. Extensive experiments across multiple reasoning benchmarks and LLM architectures show that SeRL consistently outperforms strong baselines and matches the performance of methods trained with large-scale, high-quality labeled data.

### v2

@@ 4.2 Experimental Results
Our approach matches the performance of the method using extensive data with verifiable rewards. As shown in Tab. 1, after just the first round of unsupervised RL training on generated data, our model achieves performance comparable to RL-GT. After the second round, LLaMA-3.2-3B-Instruct even outperforms RL-GT across nearly all evaluation benchmarks. LLaMA-3.2-3B-Instruct continues to improve in the third round. After three rounds of iterative training, Qwen-2.5-7B-Instruct also outperforms RL-GT on the MATH-500, ASDiv, and TabMWP benchmarks. Additional results of our method for more iterations are provided in Appendix D.2.
Our method outperforms other multi-round iterative approaches. Across all iterations, our method consistently outperforms both SR-DPO and I-RPO on LLaMA-3.2-3B-Instruct. We observe that both SR-DPO and I-RPO exhibit noticeable model degradation over multiple iterations. For SR-DPO, it performs poorly on LLaMA-3.2-3B-Instruct, while showing slightly better results on Qwen-2.5-7B-Instruct. This suggests that model-based reward mechanisms may depend on model scale, as larger models are more likely to produce accurate reward estimates. To better understand the limited effectiveness of SR-DPO, we further analyze its reward estimation accuracy in Section 4.3 (c). In contrast, I-RPO shows a decline in performance over multiple training rounds on both models. We hypothesize that this degradation results from the inclusion of a negative log-likelihood (NLL) loss term in its training objective, which may cause the model to overfit the chosen responses. As a consequence, this overoptimization reduces generalization and ultimately leads to performance drops.
Our method generalizes well to general reasoning tasks. To better evaluate generalization, we assess our method on the MMLU-Pro benchmark. As shown in Tab. 2, LLaMA-3.2-3B-Instruct achieves consistent gains over multiple iterations in certain categories, whereas Qwen-2.5-7B-Instruct shows minimal improvement. We attribute this to Qwen-2.5 already being extensively fine-tuned on MATH data, as further rule-based RL training on the MATH training set also fails to improve its MMLU-Pro performance. LLaMA-3.2-3B-Instruct shows consistent improvement in the STEM category, with additional gains in Humanities, eventually surpassing RL-GT. Performance in the Social and Other categories remain largely unchanged. This is likely because our training data is math-focused, enhancing reasoning skills that benefit STEM tasks. The improvement in Humanities is mainly due to the Law subcategory, where many questions involve logical reasoning. As the reasoning ability of the model improves, its performance on Law tasks also increases.
@@ 4.3 Ablation Study
(a) The dual-end clipped difficulty filtering mechanism effectively prevents reward hacking. As shown in Fig. 2, when we remove the difficulty filtering, the model exhibits clear signs of reward hacking: the average reward continues to increase with training steps, but the accuracy on the MATH-500 test set drops sharply after around 100 steps.
@@
Figure 2: Training curve of LLaMA-3.2-3B-Instruct with and without the online filter.
To investigate this issue, we perform a case study (Tab. E.2) and observe that although the intermediate reasoning is often correct, the model consistently ends with the final statement: “The final answer is: $\boxed{0}$.” This behavior arises because our reward function is based on majority voting. If all $n_{vote}$ sampled responses from the LLM produce the same (albeit incorrect) answer, such as $\boxed{0}$, each response receives a high reward, despite being a wrong answer. The root cause of this phenomenon lies in the difficulty of the instruction. When an instruction is too hard, the model struggles to produce consistent answers across samples, leading to unstable or misleading reward signals. To address this, we introduce a dual-end difficulty filtering mechanism to filter out instructions with insufficient answer agreement, i.e., those with $r_{\text{mean}}<\gamma_{\text{diff}}$. Similarly, overly easy instructions do not contribute meaningfully to learning and are also filtered by applying the upper threshold $r_{\text{mean}}>\gamma_{\text{easy}}$. With this dual-end clipped difficulty filtering mechanism, the reward hacking issue is mitigated in subsequent experiments.
(b) All of our proposed data filtering strategies demonstrate their effectiveness. To demonstrate the effectiveness of each filtering strategy in the online filtering process, we conducted an ablation study as shown in Tab. 3. The results show that every filtering strategy contributes to the model’s performance, as removing any of them leads to a decline in overall results. In particular, omitting the difficulty filter may cause reward hacking, as illustrated in Fig. 2.
Table 3: Pass@1 comparison under the same number of training steps with different filtering strategies ablated, using LLaMA-3.2-3B-Instruct.
Methods MATH-500 MATH-Hard ASDiv College Math TabMWP
SeRL 52.6 23.7 89.0 37.7 70.6
SeRL w/o Length Filter 47.6 23.2 87.4 36.2 60.1
SeRL w/o Keywords Filter 48.0 22.1 87.4 36.3 62.5
SeRL w/o Similarity Filter 48.8 23.3 87.3 35.6 64.6
SeRL w/o Difficulty Filter 11.6 5.1 1.5 14.0 10.2
(c) Our method effectively mitigates the bias inherently introduced by self-instruction and self-rewarding. Our proposed method is specifically designed to minimize this bias as much as possible, and matches the performance of methods using extensive data with verifiable rewards, despite the limited amount of seed data used in our method.
For the self-instruction module, we introduce an online filtering strategy that maintains data quality while promoting diversity in the generated samples. We have provided a comprehensive analysis of the generated data in terms of quality, difficulty, and diversity, as detailed in Appendix F. These results demonstrate the reliability of our self-instruction method for data generation.
@@
For the self-rewarding module, as stated in the related works [81, 87, 92], estimation bias is inevitable in self-rewarding settings. Our majority-voting reward method assigns rewards based on the consistency among multiple sampled responses, offering greater stability than scoring individual responses. To demonstrate this, we calculate the cosine similarity between the estimated rewards and the ground-truth rewards, with higher similarity indicating greater accuracy in reward estimation. As shown in Fig. 3, on both LLaMA-3.2-3B-Instruct and Qwen-2.5-7B-Instruct, the majority-voting reward demonstrates significantly higher alignment with the rule-based reward, indicating superior accuracy in reward estimation. In contrast, the low accuracy of the model-based reward on LLaMA-3.2-3B-Instruct likely accounts for the underperformance of SR-DPO on this model. Additionally, the model-based reward shows higher consistency with the rule-based reward on Qwen-2.5-7B-Instruct, which accounts for the relatively better performance of SR-DPO on Qwen-2.5-7B-Instruct. More comparisons of reward methods and alignment metric results can be found in Appendix E.1.
@@
Figure 3: Cosine similarity between rule-based reward and majority-voting / model-based reward on MATH-500. Each point represents the similarity over 16 sampled responses per instruction. "sim(GT, Model-based)" refers to the cosine similarity between the rule-based reward and the model-based reward, while "sim(GT, Majority-voting)" refers to the cosine similarity between the rule-based reward and our majority-voting reward. The dashed lines in the figure indicate the average values.
To further assess the effectiveness of our majority-voting reward, we conduct an ablation study shown in Tab. 4. In this setting, we remove the self-instruction mechanism and perform one round of majority-voting reward RL using the full MATH training set, denoted as RL-MV. Results show that this variant achieves performance comparable to RL-GT across all test datasets. Notably, LLaMA-3.2-3B-Instruct outperforms RL-GT on MATH-500, MATH-Hard, College Math, and TabMWP, while Qwen-2.5-7B-Instruct consistently surpasses RL-GT on nearly all benchmarks.
@@
(d) Our method is applicable to different RL algorithms. To verify the robustness of our method across different RL algorithms, we evaluate it under various RL setups. RLOO [1] and Reinforce++ [19] use $k_{1}$ KL divergence estimation, while GRPO [54] adopts $k_{3}$ KL divergence estimation. The formulations for both $k_{1}$ and $k_{3}$ KL divergence estimations are detailed in Appendix B. Since $k_{1}$ is integrated into the per-token reward term, we set a smaller KL coefficient for training RLOO and Reinforce++ to avoid strong KL regularization. All other hyperparameter settings remain the same. As shown in Tab. 5, all algorithms achieve comparable performance, demonstrating the generality of our approach. Among them, the Reinforce++ algorithm delivers the best overall results. The GRPO algorithm achieves relatively strong performance in the first iteration but shows continued improvement only on MATH-Hard, ASDiv, and TabMWP in subsequent rounds, while its performance declines on MATH-500 and College Math. Similarly, RLOO exhibits performance degradation on MATH-Hard and College Math after the third iteration.
Reinforce++ demonstrates greater robustness compared to GRPO and RLOO due to the following reasons: (1) GRPO normalizes the estimated advantage over the $n_{\text{vote}}$ responses for each question, whereas Reinforce++ estimates the advantage over all responses across all questions, resulting in a larger group and more stable estimation. In addition, GRPO uses an external $k_{3}$ KL estimation, which involves an exponential term. This may cause large spikes in the gradient, leading to instability during training. (2) RLOO applies a preprocessing step that subtracts the average reward of other responses from the reward of each individual response, thereby increasing the reward gap between different responses. However, since self-rewarding may introduce inaccurate reward estimations, the amplified reward differences in RLOO could lead to greater bias when the estimation is incorrect. As a result, its performance tends to be worse than that of Reinforce++.
Table 5: Comparison of SeRL performance under different reinforcement learning algorithms. Bold indicates the best performance.
@@ 5 Related Work
Self-Instruction Methods. Recent studies show that RL can significantly enhance the reasoning abilities of LLMs. However, it depends on high-quality instructions and verifiable labels, which are scarce in specialized domains such as clinical diagnostics [58, 44, 60] and aerospace engineering [36, 5, 75]. This challenge has spurred interest in synthetic data generation. Previous work has employed powerful expert models like GPT-4 to generate synthetic data [79, 13, 32, 30, 82], but these methods are costly and rely on external proprietary systems. To mitigate this dependency, self-instruction frameworks such as Wang et al. [63] bootstrap data from a small set of seed examples via few-shot generation and heuristic filtering to ensure quality and diversity. Many other studies [88, 10, 27, 11, 17] have extended the self-instruction framework, focusing on improving data efficiency, diversity, and quality. Some studies train dedicated models for instruction generation. For instance, WizardLM [74] evolves instruction data in terms of complexity, diversity, and quality, while TeaMs-RL [15] extends it by defining prompt-based augmentation actions and training an Instructor LLM via RL to apply them. In contrast, we find that existing models already possess strong instruction-generation abilities. Without training a separate model or designing additional augmentation prompts, our method leverages few-shot generation and an online filtering strategy to produce diverse, high-quality instructions (see Appendix F). Unlike these offline SFT approaches with fixed data, RL enables online data generation that adapts to the model’s evolving capability. However, it still requires verifiable labels and mechanisms to prevent reward hacking.
@@
Self-Rewarding Methods. Recent successes of RL in LLMs largely rely on large-scale instructions with verifiable rewards [73, 23, 66]. However, labeled data typically requires substantial human effort and is limited in both efficiency and quality. As a more efficient alternative, self-rewarding methods that estimate rewards without relying on verifiable labels have gained increasing attention. Model-based self-rewarding methods like Bai et al. [2] and Yuan et al. [81] use predefined rules to generate preference pairs for Direct Preference Optimization (DPO) training. Fränken et al. [9] maximize mutual information between principles and responses, while Zhang et al. [87] is based on semantic entropy clustering. RLSC [33] optimizes the model by directly maximizing its most probable response, while our majority-voting method is based on multiple responses, leading to greater stability and more consistent outputs. TTRL [92], as a concurrent work, also employs a majority-voting self-rewarding mechanism in an unsupervised reinforcement learning setting. However, our approach differs in two important aspects: (1) we generate diverse synthetic training instructions from limited data through self-instruction, unlike test-time training of TTRL on fixed test sets; (2) our method supports multiple rounds of iterative updates, enabling continual self-improvement and refinement of the policy, while TTRL only performs a single-round adaptation on a static test set.
@@ 6 Discussion
Ceiling of Self-Iteration. We begin by discussing the performance upper bound of LLMs within a single iteration. When performing RL training using only majority-voting reward, the maximum achievable Pass@1 on the training set is inherently limited by the Maj@K before training. If the LLM has strong generalization capabilities, the Pass@1 upper bound on an in-domain test set should also align with the Maj@N of the model on the test set before training. So, what is the upper bound of LLM capability across multiple iterations? If the Maj@N of the model continues to improve after each round of training, then each iteration defines a new, higher ceiling. In this case, the LLM is capable of approaching increasingly better performance bounds, suggesting the potential for unbounded, continual improvement through iterative self-training. However, as shown in Fig. 7, the Maj@16 performance of both LLaMA-3.2-3B-Instruct and Qwen-2.5-7B-Instruct on MATH-500 does not continue to improve with more iterations. This observation aligns with the analysis by Yue et al. [83], which suggests that RL with verifiable reward primarily transforms Pass@K ability into Pass@1, while the upper bound of a model is constrained by its underlying capability. Nevertheless, our method remains valuable, as it enables efficient reasoning in data-scarce scenarios through unsupervised RL training with data augmentation.
Limitation of Majority-voting Reward. In scenarios where there is no deterministic final answer such as writing tasks, or where the correctness of the process is more important such as mathematical proofs, our majority voting strategy may not be applicable. However, for most knowledge-based scenarios, even when a numeric answer is not required as in mathematical problem solving, the question can often be reformulated into a multiple-choice format, making our majority voting approach still applicable. Another limitation is that the reliability of the reward signal in self-rewarding methods can be compromised for challenging problems, as also discussed in prior work [81, 87, 92]. To mitigate this issue, we have designed a difficulty filtering strategy that removes instructions for which the model shows uncertainty in generating consistent answers. This acts as a form of curriculum learning, where the model starts training on simpler problems and gradually moves toward more complex ones. As a result, the estimation bias induced by unreliable self-rewards is reduced during training. While this strategy does not completely resolve the issue, we believe it is a promising direction and worth further exploration in future work.
@@ 7 Conclusion
In this work, we introduce SeRL, a framework designed to bootstrap LLM training from limited initial data. SeRL consists of two key components: a self-instruction module and a self-rewarding module. The self-instruction module expands the initial data by generating new instructions, with robust and effective online filtering applied to ensure quality, diversity, and appropriate difficulty. The self-rewarding module uses a majority-voting strategy to estimate rewards for generated responses, removing the need for external labels. Based on the generated data, SeRL performs standard RL training in an iterative, self-improving manner. Extensive experiments across multiple reasoning benchmarks and LLM architectures show that SeRL consistently outperforms strong baselines and matches the performance of methods trained with large-scale, high-quality labeled data.
Acknowledgement
This work is supported in part by the Hangzhou Joint Funds of the Zhejiang Provincial Natural Science Foundation of China under Grant No. LHZSD24F020001, in part by the Zhejiang Province High-Level Talents Special Support Program “Leading Talent of Technological Innovation of Ten-Thousands Talents Program” under Grant No. 2022R52046, in part by the Fundamental Research Funds for the Central Universities under Grant No. 2021FZZX001-23, in part by the National Natural Science Foundation of China (62506330), and in part by the advanced computing resources provided by the Supercomputing Center of Hangzhou City University. This research is supported by the RIE2025 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) (Award I2301E0026), administered by A*STAR, as well as supported by Alibaba Group and NTU Singapore through Alibaba-NTU Global e-Sustainability CorpLab (ANGEL).
Appendix